Find the tenth term of the geometric sequence with
a = 5 and r = 2.

Answer by stanbon(75887)

Find the tenth term of the geometric sequence with
a = 5 and r = 2.
------------
a(n) = a(1)*r^(n-1)
---
a(10) = 5*2^9 = 2560
========================
Cheers,
Stan H.
========================